2. Which of the following statements is (are) correct?
(x) Suppose that the CPI was 200 in 2006 and was 160 in 1996. Then according to the CPI, $100 in 2006 purchased the same amount of goods and services as $80 in 1996.
(y) A worker received $5 for a daily wage in 1930. If the CPI was 17.4 in 1930 and is 215.6 today then the value of that 1930 wage in today’s dollars is more than $68.25 but less than $69.75.
(z) A business firm has an agreement with its workers to index completely the wage of its employees using the CPI. The company currently pays its production line workers $12.00 an hour and is scheduled to index their wages today. If the CPI is currently 180 and was 160 a year ago, the firm should increase the hourly wages of its workers by $1.50.
A. (x), (y) and (z)
B. (x) and (y) only
C. (x) and (z) only
D. (y) and (z) only
E. (z) only


3. Which of the following statements is (are) correct?
(x) You know that a candy bar costs $0.75 today. You also know the CPI for 1962 and the CPI for today. You would use [$0.75 x (1962 CPI / today's CPI) to compute the price of the candy bar in 1962 prices.
(y) George is a service manager at a car dealership. In year 2000, he earned a salary of $38,500. Suppose the CPI is 168.1 in year 2000 and 204.6 in 2010. If George does not want to experience a reduction in his real income for that ten-year period then his nominal income in 2010 must be at least $46,860.
(z) Marvin is offered a job in Raleigh, North Carolina where the CPI is 120 and a job in New York where the CPI is 150. His job offer in Raleigh is for $54,000. The New York job needs to pay $67,500 in order to make the two salaries have approximately the same purchasing power.
A. (x), (y) and (z)
B. (x) and (y) only
C. (x) and (z) only
D. (y) and (z) only
E. (x) only
Answer #1

(2) (C)

Purchasing power of $80 in 1996, in year 2006 = $80 x (200/160) = $100 [x is right]

Value of 1930 wage in today's dollars = $5 x (215.6/17.4) = $61.95 [y is wrong]

Current wage = $12 x (180/160) = $13.5 [Increase = $13.5 - $12 = $1.5] [z is right]

(3) (A)

Price in year N prices = Current price x (CPI in year N / Current year CPI) [x is right]

Income in 2010 = $38,500 x (204.6/168.1) = $46,860 [y is right]

Minimum salary in New York = $54,000 x (150/120) = $67,500 [z is right]